A straight line of charge has length 40.00 cm and
Answer #1


Q = 12 pC, d = 40 cm , r= 15 cm

E = lamda/(2pi*eo*x)

E = 2*k*lamda/x

linear charge dencity = lamda = Q/d

E = 2*9*10^9*12*10^-12/(0.4*0.15)

E= 3.6 N/C